# Past, Present, and Future of Bug Tracking in the Generative AI Era

### v1

@@ 3.1. Overview
Once assigned, the agents localize the bug and generate a candidate patch. The developer then reviews this patch. If the developer decides not to merge, the agents repeat the localization and patch generation process until a valid solution is produced or a predefined threshold is reached. When the developer accepts the patch, the verification phase is carried out by LLM/AI agents, which attempt to verify the fix within a specified threshold of iterations. If verification by the agent fails, the responsibility shifts back to the developer, who manually fixes the bug. This manual fix is then reviewed by another developer acting as a reviewer. If the reviewer rejects the manual patch, the developer revises it until approval is achieved. The cycle resumes with the tester once the reviewer merges the patch. If the agents cannot again verify the bug within the threshold, the developer continues revising until successful verification is achieved.
When agent-based verification (or subsequent developer revisions) confirms the fix, the patch is deployed, and the user is once more asked for final verification. If the user accepts the fix, the lifecycle is completed. If the user rejects it, the process restarts from the beginning, ensuring continuous accountability, resolution, and user satisfaction. The details of each phase in this lifecycle are elaborated in the following subsections, where we describe the design, methodology, and role of agents at every stage of the process.
@@ 3.1.1. Bug Report Creation
At the same time, the research community has begun to investigate new ways to facilitate bug report creation by reducing the burden on reporters and capturing more useful information from the outset. Several studies have highlighted that critical elements such as observed behavior (OB), expected behavior (EB), and S2R are often missing or only partially provided in bug reports, particularly for web applications (Wang et al., 2025a). To address these gaps, interactive systems have been proposed that actively guide users during the reporting process. For example, BugListener (Shi et al., 2022) demonstrated how informal conversations in collaborative live chats can be automatically transformed into structured bug reports, synthesizing OBs and EBs as well as S2R. Similarly, other research has shown that automatic follow-up questions can be selected when reports are incomplete, ensuring that reporters provide the information most useful for later reproduction (Imran et al., 2021).
@@
In our framework, we extend this trajectory by employing an LLM-powered chatbot as the entry point for bug report creation. Unlike static forms, our chatbot accepts natural language input from users and adaptively engages them with follow-up questions whenever essential details are missing. This direct interaction helps resolve the asynchronous nature of current bug reporting, where users often wait hours or days for feedback, by providing immediate responses and clarification in real time. For example, if a user reports, “the login page freezes after I click submit,” the chatbot may follow up by asking, “What did you expect to happen after clicking submit?” and receive the response, “I expected to be redirected to my account dashboard.” It can then request further context, such as the browser used or whether any error message appeared. At the same time, certain environmental details such as OS, app version, or device type can often be inferred automatically from metadata without needing to ask the user directly. By the end of this short exchange, the chatbot compiles a structured report that contains the OB (page freeze), EB (dashboard redirection), and key environmental details (browser, OS, error messages), ensuring that the report is both complete and minimally burdensome for the user. This ensures that the key components of a bug report are systematically captured during the initial interaction. By grounding bug report creation in a conversational process, we aim to both reduce the cognitive effort for end-users and increase the completeness and usefulness of the reports passed downstream to subsequent phases of our framework.
@@ 3.1.2. Bug Report Enhancement
Recent studies have leveraged LLM capabilities to enhance bug report quality. LLM-driven approaches can semantically analyze bug reports and detect missing fields (Mahmud et al., 2025). They can suggest enhancements for incomplete descriptions (Bo et al., 2024) or ambiguous content (Feng and Chen, 2024). Critical information such as system state, error messages, and execution logs can be added by the agent (Wang et al., 2024e). The agent can also ensure S2R steps are complete (Emre Furkan Akyol and Tuzun, 2025). Furthermore, LLMs maintain consistency with prior examples and best practices in bug reporting (Soltani et al., 2020). This process ensures that enhanced reports are actionable for developers (Song et al., 2023).
In our methodology, the bug report enhancement process adopts this LLM-driven approach. The agent first analyzes the submitted report and evaluates its quality against criteria such as completeness, clarity, and conciseness that are defined by Bettenburg et al. (Bettenburg et al., 2007). Based on this assessment, the agent proposes enhancements, filling missing details or rephrasing ambiguous descriptions (Mahmud et al., 2025; Emre Furkan Akyol and Tuzun, 2025). Both the original and enhanced versions of the bug report are stored in the database, allowing developers or stakeholders to reference the original submission if necessary. For all downstream tasks in our proposed bug tracking workflow, such as bug reproduction, localization, and no-code fixes or patch generation, the enhanced version is used as input, as well as the original report, for developers’ later references.
@@ 3.1.3. Bug Reproduction
In early systems, bug reproduction was manual: test engineers, customer support, or developers attempted to follow the S2R included in the report, often needing to contact users for clarification when the description was incomplete or ambiguous. Research gradually moved toward automation. Techniques emerged to extract reproduction steps from natural language descriptions (Zhao et al., 2019a) or to leverage runtime information such as logs, execution traces, or telemetry (Wang et al., 2024a). In mobile and GUI domains, specialized approaches replayed user reviews, event sequences, or monitored app interactions to recreate failures (Li et al., [n. d.]). These works highlight how the community attempted to reduce reliance on manual effort, though success was often limited to narrow domains or specific application types.
@@
With the rise of LLMs, bug reproduction has increasingly been reframed as a generative task. LLMs can read bug descriptions and produce candidate test cases or executable scenarios that explicitly trigger the reported failure, thereby supporting the reproduction of the underlying bug. LIBRO (Kang et al., 2023) demonstrated that LLMs are capable of generating test cases from natural language bug reports. Other studies integrated visual context: ReCDroid+ (Zhao et al., 2019b) and vision-based reproduction methods showed how screenshots and GUI interactions can be leveraged to systematically replay UI-intensive failures (Wang et al., 2025d). More recently, feedback-driven methods like ReBL (Wang et al., 2024e) explored iterative prompting and adaptation, where an LLM interacts with execution feedback to refine its reproduction strategy over multiple rounds. BugCraft (Yapağcı et al., 2025) extended this idea into the game domain, showing how iterative synthesis and refinement can be applied to highly interactive environments. Most recently, work at Google has shown that agentic BRT (Bug Reproduction Test) generation, where an LLM agent produces fail-to-pass tests that both reproduce and validate bugs, can significantly improve industrial-scale automated program repair by providing stronger debugging and validation signals (Cheng et al., 2025).
@@
Building on this trajectory, our methodology operationalizes bug reproduction as an iterative loop tightly coupled with bug report enhancement. Once a bug report is enhanced, the agents attempt automated reproduction in a controlled sandboxed or containerized environment, ensuring determinism and isolation from external noise (Chen et al., 2020). For GUI bugs, this involves simulating user interactions based on screenshots or videos (Liu et al., 2024; Zhao et al., 2022), while backend and API bugs may be reproduced by replaying logs or workloads reconstructed from user sessions (Wang et al., 2024a). If reproduction succeeds, the system generates an executable artifact that can be passed to verification and integrated into regression pipelines.
@@
If reproduction fails, the system increments an iteration counter and loops back to the bug report enhancement phase. Here, LLM agents refine the S2R by reordering actions, adjusting parameters, adding missing environmental constraints, or even requesting specific artifacts such as logs or screenshots (Erfani Joorabchi et al., 2014; Kang et al., 2024b; Feng and Chen, 2024). The refined report is then re-executed, and this loop continues until reproduction succeeds or a threshold is reached. At that point, the case escalates to customer support for manual handling.
@@
This iterative loop transforms reproduction from a one-shot attempt into a self-correcting process, where each failure strengthens the next attempt. Bug reports thus become dynamic artifacts that evolve through agentic refinement until a reproducible scenario is achieved. By feeding these successful reproductions into regression test suites, our framework not only reduces TTR but also creates an expanding safety net to prevent regressions in future builds (Mahmud et al., 2025).
@@ 3.1.4. Bug Classification
Once a bug is successfully reproduced, the next step is classification. Traditional bug triaging typically requires substantial manual intervention to evaluate factors such as severity, impact, and priority. Practitioners look for cues in the bug report, such as affected features, system logs, and error messages, to gauge how urgently a fix is needed. This manual process can be time-consuming and prone to human error, as it often depends on individual judgment and expertise.
@@ 3.1.5. Bug-Feature Traceability
Automating this process minimizes the risk of human error and overlooked relationships that commonly occur in manual triage. It also facilitates cross-team communication by providing a shared, up-to-date view of the relationship between bugs and features. Ultimately, maintaining accurate bug–feature trace links supports a more streamlined and data-driven defect resolution process, helping to reduce maintenance costs and improve overall product quality.
@@ 3.1.6. Bug Validity Check
Building on these studies, our proposed workflow employs an LLM agent to perform the validity check by analyzing S2R, logs, and error messages while also handling noise and ambiguity in real-world reports. The LLM agent can also use traditional ML models to make a decision about the bug report’s validity. To further strengthen its decision-making, the agent cross-references historical bug reports and project documentation in a retrieval-augmented generation (RAG) setting as in Dinç et al.’s study (Dinç and Tüzün, 2025).
@@ 3.1.7. Bug Assigner
With the advent of deep learning and LLMs, a new generation of approaches emerged. Transformer-based methods such as BERT (Devlin et al., 2019) have been applied to bug triage, enabling richer semantic representations of bug reports and improving accuracy (Lee et al., 2023). Ensemble approaches that combine multiple language models have further improved robustness and prediction performance (Kumar Dipongkor, 2024). Beyond software maintenance, similar ideas have been explored in incident management: COMET demonstrated that LLMs can provide accurate and interpretable assignment of cloud incidents at Microsoft, showing that LLMs can reliably map natural language reports to the right resolver (Wang et al., 2024b).
@@ 3.1.8. Bug Handling with No-Code Fixes
If a bug report is deemed invalid, it often requires no-code fixes. No-code fixes resolve such issues by making changes outside of the application’s source code, typically involving adjustments to configuration files, server settings, user permissions, or system-level parameters. In current bug tracking platforms, these fixes are usually proposed and applied by customer support staff rather than PMs or developers, since they often relate to usage errors, misconfigurations, or environment-specific issues rather than programming defects.
@@ 3.1.9. Bug Localization
Once a PM approves a bug for fixing and confirms the assignment, the process of bug localization begins. Historically, this has been a challenging task, as developers must manually sift through vast and complex codebases to find the exact line of code responsible for an issue (Ghosh and Singh, 2019). Traditional bug localization techniques, such as Information Retrieval-Based Fault Localization (Xia and Lo, 2023) and Spectrum-Based Fault Localization (Wen et al., 2019), have been effective, but they are often complemented by other methods. Slicing-based fault localization (Wong et al., 2023) narrows down the search space by identifying relevant parts of the code that could influence the program’s behavior, while Mutation-based fault localization (Ghanbari et al., 2023) generates a set of mutants (tiny, seeded bugs) to see if they can be killed by the failing test cases, thus helping to pinpoint the real bug. Still, these methods often rely on statistical analysis of test coverage rather than a deep understanding of the code’s semantic meaning.
The recent integration of LLMs has introduced a new paradigm, significantly enhancing debugging efficiency by going beyond traditional methods. Researchers have proposed various frameworks and systems to leverage LLMs for localization (Yang et al., 2024; Kang et al., 2024a; Do Viet and Markov, 2023; Li et al., 2025). For instance, the FlexFL framework (Xu et al., 2025) introduces a two-stage approach. The first stage uses traditional fault localization techniques to narrow down potential bug locations, and the second stage uses an LLM-based agent to perform a deeper analysis of the reduced search space. Similarly, SOAPFL (Qin et al., 2025) formalizes the bug localization process into a Standard Operating Procedure (SOP) that mimics a human developer’s workflow, comprising comprehension, navigation, and confirmation steps. These studies showcase how LLMs can directly understand the semantic meaning of the code itself, enabling a more intuitive and accurate approach to finding the root cause.
@@ 3.1.10. Patch Generation
The recent advancements in AI-driven patch generation have revolutionized this part of the bug-fixing process, and we would utilize existing studies to establish our semi-automated bug-tracking system. This is a well-studied field, with a growing body of research demonstrating how LLMs can automatically generate code fixes for a wide range of bugs (Xia et al., 2023; Huang et al., 2024; Kang et al., 2025; Wang et al., 2025c).
@@
By using this approach, instead of manually sifting through the code to create a patch, developers can review and refine AI-generated suggestions, freeing them to focus on more complex, creative tasks. By minimizing human error and proactively suggesting optimal fixes, the system ensures that the final product is more robust and secure.
3.1.11. Patch Verification
In the patch verification step, our system assesses the correctness of generated patches and iteratively refines them to ensure reliability and semantic correctness. Foundational work in APR highlighted the importance of rigorous verification to move candidate patches from plausibility to practical adoption (Qi et al., 2014; Xiong et al., 2017).
Recent studies have explored leveraging LLMs for patch generation and validation as well as verification. LLM-driven approaches can generate test cases to verify candidate patches (Li and Yuan, 2024), detect semantic deviations (Wang et al., 2024d), and iteratively refine patches based on verification feedback (Le-Cong et al., 2023). These methods demonstrate the potential of LLMs to improve patch correctness, coverage, and production readiness.
Building on these insights, our methodology integrates an LLM agent into the patch verification workflow. After a patch is generated, the agent validates it against both developer-provided and automatically synthesized oracles, utilizing artifacts generated in reproduction as detailed in Section 3.1.3. Regression tests are executed automatically to confirm that the bug is resolved without introducing new failures. In this HIL setting, test engineers supervise the verification process, reviewing the results and ensuring that candidate patches meet quality standards. Until a predefined number of iterations threshold, our proposed workflow goes back to bug localization and generates new patches in cases of verification or regression testing failures. This approach bridges the gap between APR and practical patch adoption, moving beyond syntactic plausibility toward semantically correct, production-ready fixes. If the agent cannot verify within a specified number of iterations, the verification task would be delegated to the test engineer.

### v2

@@ 3.1. Overview
Once assigned, the agents localize the bug and generate a candidate patch. The developer then reviews this patch. If the developer decides not to approve the generated patch, the agents repeat the localization and patch generation process until a valid solution is produced or a predefined threshold is reached. When the developer approves the patch, the verification phase is carried out by LLM/AI agents, which attempt to verify the fix within a specified threshold of iterations. If verification by the agent fails, the responsibility shifts back to the developer, who manually fixes the bug. This manual fix is then reviewed by another developer acting as a reviewer. If the reviewer rejects the manual patch, the developer revises it until approval is achieved. The cycle resumes with the tester once the reviewer approves the patch. If the agents cannot again verify the bug within the threshold, the developer continues revising until successful verification is achieved.
When agent-based verification (or subsequent developer revisions) confirms the fix, the patch is deployed, and the user is once more asked for final verification. If the user accepts the fix, the lifecycle is completed. If the user rejects it, the process restarts from the beginning, ensuring continuous accountability, resolution, and user satisfaction. Finally, we note that duplicate bug detection is intentionally not modeled as a standalone stage in this lifecycle. Rather than representing a single linear step, duplicate detection functions as a cross-cutting capability that can be applied at multiple points in the workflow, including during initial report creation, reproduction, classification, and validation. Because its role is inherently supportive of several phases rather than constituting a distinct phase of its own, we integrate it implicitly throughout the pipeline instead of depicting it as a separate stage. The details of each phase in this lifecycle are elaborated in the following subsections, where we describe the design, methodology, and role of agents at every stage of the process.
@@ 3.1.1. Bug Report Creation
At the same time, the research community has begun to investigate new ways to facilitate bug report creation by reducing the burden on reporters and capturing more useful information from the outset. Several studies have highlighted that critical elements such as observed behavior (OB), expected behavior (EB), and S2R are often missing or only partially provided in bug reports, particularly for web applications (Wang et al., 2025b). To address these gaps, interactive systems have been proposed that actively guide users during the reporting process. For example, BugListener (Shi et al., 2022) demonstrated how informal conversations in collaborative live chats can be automatically transformed into structured bug reports, synthesizing OBs and EBs as well as S2R. Similarly, other research has shown that automatic follow-up questions can be selected when reports are incomplete, ensuring that reporters provide the information most useful for later reproduction (Imran et al., 2021).
@@
In our framework, we extend this trajectory by employing an LLM-powered chatbot as the entry point for bug report creation. Unlike static forms, our chatbot accepts natural language input from users and adaptively engages them with follow-up questions whenever essential details are missing. This direct interaction helps resolve the asynchronous nature of traditional bug reporting, where users often wait hours or days for feedback, by providing immediate responses and clarification in real time. For example, if a user reports, “the login page freezes after I click submit,” the chatbot may follow up by asking, “What did you expect to happen after clicking submit?” and receive the response, “I expected to be redirected to my account dashboard.” It can then request further context, such as the browser used or whether any error message appeared. At the same time, certain environmental details such as OS, app version, or device type can often be inferred automatically from metadata without needing to ask the user directly. By the end of this short exchange, the chatbot compiles a structured report that contains the OB (page freeze), EB (dashboard redirection), and key environmental details (browser, OS, error messages), ensuring that the report is both complete and minimally burdensome for the user. This ensures that the key components of a bug report are systematically captured during the initial interaction. By grounding bug report creation in a conversational process, we aim to both reduce the cognitive effort for end-users and increase the completeness and usefulness of the reports passed downstream to subsequent phases of our framework.
@@ 3.1.2. Bug Report Enhancement
Recent studies have leveraged LLM capabilities to enhance bug report quality. LLM-driven approaches can semantically analyze bug reports and detect missing fields (Mahmud et al., 2025). They can suggest enhancements for incomplete descriptions (Bo et al., 2024) or ambiguous content (Feng and Chen, 2024). Critical information such as system state, error messages, and execution logs can be added by the agent (Wang et al., 2024b). The agent can also ensure S2R steps are complete (Emre Furkan Akyol and Tuzun, 2026). Furthermore, LLMs maintain consistency with prior examples and best practices in bug reporting (Soltani et al., 2020). This process ensures that enhanced reports are actionable for developers (Song et al., 2023).
In our methodology, the bug report enhancement process adopts this LLM-driven approach. The agent first analyzes the submitted report and evaluates its quality against criteria such as completeness, clarity, and conciseness that are defined by Bettenburg et al. (Bettenburg et al., 2007). Based on this assessment, the agent proposes enhancements, filling missing details or rephrasing ambiguous descriptions (Mahmud et al., 2025; Emre Furkan Akyol and Tuzun, 2026). Both the original and enhanced versions of the bug report are stored in the database, allowing developers or stakeholders to reference the original submission if necessary. For all downstream tasks in our proposed bug tracking workflow, such as bug reproduction, localization, and no-code fixes or patch generation, the enhanced version is used as input, as well as the original report, for developers’ later references.
@@ 3.1.3. Bug Reproduction
In early systems, bug reproduction was manual: test engineers, customer support, or developers attempted to follow the S2R included in the report, often needing to contact users for clarification when the description was incomplete or ambiguous. Research gradually moved toward automation. Techniques emerged to extract reproduction steps from natural language descriptions (Zhao et al., 2019a) or to leverage runtime information such as logs, execution traces, or telemetry (Wang et al., 2024a). In mobile and GUI domains, specialized approaches replayed user reviews, event sequences, or monitored app interactions to recreate failures (Li et al., ). These works highlight how the community attempted to reduce reliance on manual effort, though success was often limited to narrow domains or specific application types.
@@
With the rise of LLMs, bug reproduction has increasingly been reframed as a generative task. LLMs can read bug descriptions and produce candidate test cases or executable scenarios that explicitly trigger the reported failure, thereby supporting the reproduction of the underlying bug. LIBRO (Kang et al., 2023) demonstrated that LLMs are capable of generating test cases from natural language bug reports. Other studies integrated visual context: ReCDroid+ (Zhao et al., 2019b) and vision-based reproduction methods showed how screenshots and GUI interactions can be leveraged to systematically replay UI-intensive failures (Wang et al., 2025c). More recently, feedback-driven methods like ReBL (Wang et al., 2024b) explored iterative prompting and adaptation, where an LLM interacts with execution feedback to refine its reproduction strategy over multiple rounds. BugCraft (Yapağcı et al., 2025) extended this idea into the game domain, showing how iterative synthesis and refinement can be applied to highly interactive environments. Most recently, work at Google has shown that agentic BRT (Bug Reproduction Test) generation, where an LLM agent produces fail-to-pass tests that both reproduce and validate bugs, can significantly improve industrial-scale automated program repair by providing stronger debugging and validation signals (Cheng et al., 2025).
@@
Building on this trajectory, our methodology operationalizes bug reproduction as an iterative loop tightly coupled with bug report enhancement. Once a bug report is enhanced, the agents attempt automated reproduction in a controlled sandboxed or containerized environment, ensuring determinism and isolation from external noise such as interrupts, concurrent kernel activity, and asynchronous hardware effects (Chen et al., 2020). For GUI bugs, this involves simulating user interactions based on screenshots or videos (Liu et al., 2024; Zhao et al., 2022), while backend and API bugs may be reproduced by replaying logs or workloads reconstructed from user sessions (Wang et al., 2024a). If reproduction succeeds, the system generates an executable artifact that can be passed to verification and integrated into regression pipelines.
@@
If reproduction fails, the system increments an iteration counter and loops back to the bug report enhancement phase. In this phase, LLM agents refine the S2R by reordering actions, adjusting parameters, adding missing environmental constraints, and, following established practices in prior work, actively requesting additional artifacts such as logs or screenshots (Erfani Joorabchi et al., 2014; Kang et al., 2024b; Feng and Chen, 2024). The refined report is then re-executed, and this loop continues until reproduction succeeds or a threshold is reached. At that point, the case escalates to customer support for manual handling.
@@
This iterative loop transforms reproduction from a one-shot attempt into a self-correcting process, where each failure strengthens the next attempt. Bug reports thus become dynamic artifacts that evolve through agentic refinement until a reproducible scenario is achieved. By feeding these successful reproductions into regression test suites, our framework creates an expanding safety net to prevent regressions in future builds (Mahmud et al., 2025).
@@ 3.1.4. Bug Classification
Once a bug is successfully reproduced, the next step is classification. Traditional bug classification typically requires substantial manual intervention to evaluate factors such as severity, impact, and priority (Anvik et al., 2006). Practitioners look for cues in the bug report, such as affected features, system logs, and error messages, to gauge how urgently a fix is needed. This manual process can be time-consuming and prone to human error, as it often depends on individual judgment and expertise.
@@ 3.1.5. Bug-Feature Traceability
Automating this process minimizes the risk of human error and overlooked relationships that commonly occur in manual triage. It also facilitates cross-team communication by providing a shared, up-to-date view of the relationship between bugs and features. Ultimately, maintaining accurate bug-feature trace links supports a more streamlined and data-driven defect resolution process, helping to reduce maintenance costs and improve overall product quality.
@@ 3.1.6. Bug Validity Check
Building on these studies, our proposed workflow employs an LLM agent to perform the validity check by analyzing S2R, logs, and error messages while also handling noise and ambiguity in real-world reports. The LLM agent can also use traditional ML models to make a decision about the bug report’s validity. To further strengthen its decision-making, the agent cross-references historical bug reports and project documentation in a retrieval-augmented generation (RAG) setting as in Dinç and Tüzün’s study (Dinç and Tüzün, 2025).
@@ 3.1.7. Bug Assigner
With the advent of deep learning and LLMs, a new generation of approaches emerged. Transformer-based methods such as BERT (Devlin et al., 2019) have been applied to bug triage, enabling richer semantic representations of bug reports and improving accuracy (Lee et al., 2023). Ensemble approaches that combine multiple language models have further improved robustness and prediction performance (Kumar Dipongkor, 2024). Beyond software maintenance, similar ideas have been explored in incident management: COMET demonstrated that LLMs can provide accurate and interpretable assignment of cloud incidents at Microsoft, showing that LLMs can reliably map natural language reports to the right resolver (Wang et al., 2024f).
@@ 3.1.8. Bug Handling with No-Code Fixes
If a bug report is deemed invalid, it often requires no-code fixes. No-code fixes resolve such issues by making changes outside of the application’s source code, typically involving adjustments to configuration files, server settings, user permissions, or system-level parameters. In traditional bug tracking platforms, these fixes are usually proposed and applied by customer support staff rather than PMs or developers, since they often relate to usage errors, misconfigurations, or environment-specific issues rather than programming defects.
@@ 3.1.9. Bug Localization
Once a PM approves a bug for fixing and confirms the assignment, the process of bug localization begins. Historically, this has been a challenging task, as developers must manually sift through vast and complex codebases to find the exact line of code or code chunk responsible for an issue (Ghosh and Singh, 2019). Traditional bug localization techniques, such as Information Retrieval-Based Fault Localization (Xia and Lo, 2023) and Spectrum-Based Fault Localization (Wen et al., 2019), have been effective, but they are often complemented by other methods. Slicing-based fault localization (Wong et al., 2023) narrows down the search space by identifying relevant parts of the code that could influence the program’s behavior, while Mutation-based fault localization (Ghanbari et al., 2023) generates a set of mutants (tiny, seeded bugs) to see if they can be killed by the failing test cases, thus helping to pinpoint the real bug. Still, these methods often rely on statistical analysis of test coverage rather than a deep understanding of the code’s semantic meaning.
The recent integration of LLMs has introduced a new paradigm, significantly enhancing debugging efficiency by going beyond traditional methods. Researchers have proposed various frameworks and systems to leverage LLMs for localization (Yang et al., 2024; Kang et al., 2024a; Do Viet and Markov, 2023; Li et al., 2025). For instance, the FlexFL framework (Xu et al., 2025) introduces a two-stage approach. The first stage uses traditional fault localization techniques to narrow down potential bug locations, and the second stage uses an LLM-based agent to perform a deeper analysis of the reduced search space. Similarly, SOAPFL (Qin et al., 2025) formalizes the bug localization process into a procedure that mimics a human developer’s workflow, comprising comprehension, navigation, and confirmation steps. These studies showcase how LLMs can directly understand the semantic meaning of the code itself, enabling a more intuitive and accurate approach to finding the root cause.
@@ 3.1.10. Patch Generation
The recent advancements in AI-driven patch generation have revolutionized this part of the bug-fixing process, and we would utilize existing studies to establish our semi-automated bug tracking system. This is a well-studied field, with a growing body of research demonstrating how LLMs can automatically generate code fixes for a wide range of bugs (Xia et al., 2023; Huang et al., 2024a; Kang et al., 2025; Wang et al., 2025d).
@@
By using this approach, instead of manually editing the code, developers can review and refine AI-generated suggestions, freeing them to focus on more complex, creative tasks. By minimizing human error and proactively suggesting optimal fixes, the system ensures that the final product is more robust and secure.
3.1.11. Bug Fix Verification
In the bug fix verification step, our system assesses the correctness of generated patches and iteratively refines them to ensure reliability and semantic correctness. Foundational work in APR highlighted the importance of rigorous verification to move candidate patches from plausibility to practical adoption (Qi et al., 2014; Xiong et al., 2017).
Recent studies have explored leveraging LLMs for patch generation and validation, as well as verification. LLM-driven approaches can generate test cases to verify candidate patches (Li and Yuan, 2024), detect semantic deviations (Wang et al., 2024d), and iteratively refine patches based on verification feedback (Le-Cong et al., 2023). These methods demonstrate the potential of LLMs to improve patch correctness, coverage, and production readiness.
Building on these insights, our methodology integrates an LLM agent into the bug fix verification workflow. Crucially, this agent is designed not merely to act as a passive test runner, but as a rigorous evaluator that actively examines patches for deceptive or trivial fixes. To prevent the generation of trivial or deceptive fixes, such as simply removing a failing assertion or skipping a test, the system enforces strict constraints, such as existing tests cannot be deleted or modified without explicit human authorization.
Furthermore, the agent employs active test augmentation to ensure the patch is not merely overfit to the specific failure. By generating new test cases that target the modified code paths, the system verifies that the patch preserves existing functionality and does not introduce regressions. In this HIL setting, test engineers supervise the verification process, reviewing the results and ensuring that candidate patches meet quality standards. Until a predefined number of iterations threshold is reached, our proposed workflow reverts to bug localization and generates new patches in cases of verification or regression testing failures. This approach bridges the gap between APR and practical patch adoption, moving beyond syntactic plausibility toward semantically correct, production-ready fixes. If the agent cannot verify within a specified number of iterations, the verification task would be delegated to the test engineer.